OpenStudy (anonymous):

5 square root 6x +8 square root 14x -9 square root 6x + 18 square root 14x

OpenStudy (radar):

\[5\sqrt{6x} +8\sqrt{14x}-9\sqrt{6x}+18\sqrt{14x}\]Is that the problem?

OpenStudy (anonymous):

yes indeed

OpenStudy (radar):

I think I made an error

OpenStudy (anonymous):

oh, well if you combine, wouldn't it be -4 square root 6x + 26 square root 14x?

OpenStudy (anonymous):

the same answer, but 26 instead of 22

OpenStudy (anonymous):

right?

OpenStudy (radar):

Yes, for some reason my computer put me to another problem. Yes

OpenStudy (radar):

Do you have a list of options?

OpenStudy (anonymous):

yes, I just checked my list of options and there was one that matched the exact answer

OpenStudy (radar):

I ask because, there can be some more done, but I can't see any advantage to do it.

OpenStudy (radar):

Go for it brittneyy18rm

OpenStudy (anonymous):

thanks, I have a few more questions I am about to post that are similar to this one would you mind helping me through them?

OpenStudy (radar):

No I don't mind

OpenStudy (anonymous):

thanks a bunch

OpenStudy (radar):

I am sure the solutions would be similar to what we just did.